# Unraveling Zeroth-Order Optimization through the Lens of Low-Dimensional Structured Perturbations

### v1

@@ Appendix D Detailed Description of MeZO-BCD
MeZO with Block Coordinate Descent (MeZO-BCD) follows a block-wise zeroth-order optimization strategy, where the parameters are partitioned into $N$ disjoint blocks, $\mathbf{B}=\{B_{i}\}_{i=1}^{N}$. In our method, each block corresponds to all the parameters of a single Transformer decoder layer, ensuring structured and efficient updates.
D.1 Block-Wise Zeroth-Order Optimization
At each iteration, instead of updating all parameters simultaneously, MeZO-BCD selects only one block to update. The active block index $j$ is determined based on one of the following update strategies:
  • Random permutation-based selection: Instead of sampling blocks independently at random, a random permutation of $[N]$ is precomputed and followed for $N$ steps. After each full cycle, a new random permutation is generated. This ensures that all blocks are updated exactly once per cycle while maintaining randomness.
  • Flip-flop selection: The block index oscillates sequentially between $1$ and $N$, ensuring all layers are updated in a cyclic order.
  • Ascending order: Blocks are updated sequentially from $1$ to $N$, then restart from $1$.
  • Descending order: Blocks are updated sequentially from $N$ to $1$, then restart from $N$.
Among these strategies, random permutation-based and flip-flop selection generally showed better performance in most cases. The random permutation approach ensures that each block is updated exactly once per cycle while avoiding systematic biases introduced by sequential ordering. The flip-flop strategy ensures a balanced coverage of all layers over time. Ascending and descending orders, while systematic, tended to introduce biases in weight updates, increasing the likelihood of converging to a suboptimal solution.
D.2 Efficient Implementation via Block Prefix Matching
In the actual implementation, we efficiently partition the model parameters by maintaining a list of prefixes corresponding to each block. Instead of explicitly storing parameter indices, we apply perturbations only to parameters whose names match a stored prefix. Furthermore, we treat the embedding layers and the language modeling head (lm_head) as a single block, in addition to the Transformer decoder layers. This ensures that parameter updates remain consistent across all critical model components.
For example, in OPT-1.3B, the parameter prefixes used for block partitioning are:
  • "model.decoder.layers.0.", "model.decoder.layers.1.", $\dots$, "model.decoder.layers.23."
  • "model.decoder.embed_tokens." (Embedding layer)
  • "lm_head." (Language model head)
At each training step, the function PerturbParameters applies perturbations only to parameters whose names contain the prefix corresponding to the currently selected block.
D.3 Computational Benefit of Block Updates
A key advantage of MeZO-BCD is that each training step only requires generating a random matrix for a single block. Since random matrix generation is required for zeroth-order gradient estimation, performing this operation on only one block per step significantly reduces computation time.
This efficiency gain is particularly important given that zeroth-order optimization relies solely on forward passes. Unlike first-order optimization, where gradient computation dominates the computational cost, even seemingly small operations such as generating random perturbations can introduce substantial overhead in zeroth-order methods. By restricting perturbation generation to a single Transformer decoder layer per step, MeZO-BCD minimizes this overhead while maintaining low-dimensionally structured updates.
Why BCD is More Effective in Zeroth-Order Optimization.
In first-order (FO) optimization, applying block coordinate descent (BCD) can lead to significant computational and memory inefficiencies. The primary reason is that backpropagation propagates gradients through all subsequent layers, meaning that when updating a block closer to the input layer, the computational and memory cost grows substantially due to the need to compute and store gradients for all proceeding layers.
However, in zeroth-order (ZO) optimization, there is no backpropagation, and the gradient is estimated based solely on the perturbed function evaluations. This means that only the selected block is affected at each step, with no additional computational overhead from later layers. As a result, ZO-BCD fully realizes the ideal efficiency of block-wise updates, allowing purely localized updates without unnecessary computational dependencies. This makes BCD particularly well-suited for ZO, as it achieves the desired reduction in per-step complexity without the typical drawbacks encountered in FO optimization.
D.4 On the Choice of Block Partitioning Granularity
The choice of block granularity is crucial in determining both computational efficiency and optimization behavior. We considered several partitioning strategies:
  • Finer partitions: Blocks corresponding to individual nn.Linear layers, or grouping all self-attention modules separately from MLPs.
  • Coarser partitions: Merging two consecutive decoder layers into a single block.
Through empirical evaluation, we found that:
  • Finer partitions significantly slowed down convergence, likely due to excessive variance in the zeroth-order gradient estimates. The updates became too fragmented, making optimization inefficient.
  • Coarser partitions showed no major advantage over learning rate tuning, meaning that increasing the block size did not offer a meaningful improvement in optimization dynamics.
Based on these findings, we standardize one Transformer decoder layer as a single block, balancing efficiency and convergence speed.
Algorithm 1 MeZO with Block Coordinate Descent (MeZO-BCD) - Random Order
@@ D.4 On the Choice of Block Partitioning Granularity
0:  parameters $\mathbf{\theta}\in\mathbb{R}^{d}$, loss function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule $\{\eta_{t}\}_{t=1}^{T}$, block partitioning $\mathbf{B}=\{B_{i}\}_{i=1}^{N}$
0:
  Initialize block permutation $\pi\leftarrow\texttt{Shuffle}([N])$
  for $t=1,\ldots,T$ do
     Sample minibatch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
     Update current active block index $j\leftarrow\pi[t\bmod N]$
     $\ell_{+}\leftarrow\mathcal{L}(\mathbf{\theta};\mathcal{B})$
     $\ell_{-}\leftarrow\mathcal{L}(\mathbf{\theta};\mathcal{B})$
     projected_grad $\leftarrow(\ell_{+}-\ell_{-})/2\mu$
     Reset random number generator with seed $s$
     for $\theta_{i}\in\mathbf{\theta}_{B_{j}}$ do
        $u\sim\mathcal{N}(0,1)$
        $\theta_{i}\leftarrow\theta_{i}-\eta_{t}$ projected_grad * $u$
     end for
     if $t\bmod N=0$ then
        $\pi\leftarrow\texttt{Shuffle}([N])$
     end if
  end for
  function PerturbParameters($\mathbf{\theta},\mu,s$)
     Reset random number generator with seed $s$
     for $\theta_{i}\in\mathbf{\theta}$ do
        $u\sim\mathcal{N}(0,1)$
        $\theta_{i}\leftarrow\theta_{i}+\mu u$
     end for
     return $\mathbf{\theta}$
  end function
Algorithm 2 MeZO with Block Coordinate Descent (MeZO-BCD) - Deterministic Order
0:  parameters $\mathbf{\theta}\in\mathbb{R}^{d}$, loss function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule $\{\eta_{t}\}_{t=1}^{T}$, block partitioning $\mathbf{B}=\{B_{i}\}_{i=1}^{N}$
@@
     Update current active block index $j\leftarrow\mathrm{\texttt{UpdateBlockIdx}}(t,N)$
@@
  function PerturbParameters($\mathbf{\theta},\mu,s$)
@@
  end function
  function UpdateBlockIdx($t,N$)
     if ascending then
        $i\leftarrow(t\bmod N)+1$
     else if descending then
        $i\leftarrow N-(t\bmod N)$
     else if Block update order is flip-flop then
        $i\leftarrow N-1-|(t\bmod(2N-2))-(N-1)|$
@@ E.1 Randomized Quadratic Minimization
For the experiments in Section 4, the perturbation scale is globally set to $\mu=10^{-4}$ unless stated otherwise. The Hessian-like matrix $\mathbf{H}$ is generated as follows. The matrix is constructed as a block-diagonal structure, where each block is a randomly generated positive semi-definite matrix. Specifically, the $d\times d$ matrix $\mathbf{H}$ is divided into $B$ non-overlapping blocks, each of size $\frac{d}{B}\times\frac{d}{B}$. For each block, a random $\frac{d}{B}\times\frac{d}{B}$ matrix is sampled, and an orthogonal matrix $Q_{i}$ is obtained via QR decomposition. Given a predefined maximum eigenvalue $\lambda_{\max}^{(i)}$ for each block, the diagonal matrix $\Lambda^{(i)}$ is constructed as:
$\Lambda^{(i)}=\mathrm{diag}(\underbrace{\lambda_{\max}^{(i)},\ldots,0.1\times\lambda_{\max}^{(i)}}_{r},\underbrace{0,\ldots,0}_{\frac{d}{B}-r}),$
@@
$\mathbf{H}_{i}=Q_{i}\Lambda^{(i)}Q_{i}^{\top}.$
@@
This process ensures that $\mathbf{H}$ maintains a block-diagonal structure while preserving the spectral properties within each block. The following Python function is used to generate $\mathbf{H}$:
@@ E.1.1 Configuration for Figure 1
To examine the influence of the effective overlap $\rho$, we use a dense Hessian matrix (i.e., $B=1$). The effective overlap $\rho$ is controlled by introducing a parameter $\gamma\in[0,1]$, which determines the proportion of Hessian eigenvectors incorporated into the low-rank projection matrix $M$. Specifically, $\lceil s\gamma\rceil$ eigenvectors (where $s=\mathrm{srank}(M)$) are randomly selected, while the remaining vectors are randomly generated and orthogonalized against the selected ones.
Formally, let the Hessian matrix $\mathbf{H}$ have rank $r$, and let $s$ denote the stable rank of $M$. Denote the eigenvectors of $\mathbf{H}$ corresponding to nonzero eigenvalues as $\{\mathbf{q_{1}},\mathbf{q_{2}},\dots,\mathbf{q_{r}}\}$. To construct $M$ with a controlled overlap $\rho$:
@@
    Randomly selected $s\gamma$ eigenvectors are set as columns of $M_{1}$ (assuming $s\gamma$ is an integer for simplicity).
@@
    A random $d\times(1-s\gamma)$ matrix $R$ is generated.
@@
    The QR decomposition is applied to $R-MM^{\top}R$ to obtain an orthogonal matrix $M_{2}$ of size $d\times(1-s\gamma)$.
@@
    The final projection matrix is constructed as $M=[M_{1},M_{2}]$.
@@
For this experiment, we set $d=256$, the Hessian rank to 64, and the stable rank of $M$ to 64. This ensures that when $\gamma=1.0$, the projection matrix fully captures the Hessian eigenvectors. We use a learning rate of $\eta=10^{-3}$, a maximum of 1000 steps, and test five values of $\gamma$:
@@
Experimental Setup for the Right Panel of Figure 1.
@@
  • The target loss is set to the minimal loss achieved when $\gamma=0.0$.
@@
E.1.2 Configuration for Figure 2
To evaluate the effective overlap $\rho$ across different methods, we conduct randomized quadratic minimization experiments using a block-diagonal Hessian matrix with heterogeneous eigenspectra. To ensure a broader range of $\mathrm{srank}(M)$ values, the problem dimension is set to $d=1024$, with the Hessian consisting of 16 diagonal blocks, each of rank 16. To introduce heterogeneity in the eigenspectrum across blocks, the eigenvalues in each block are set by randomly choosing a reference from $\{10,40,70,100\}$ and sampling each eigenvalue as an integer within ±$\pm 2$ of it.
Measuring Effective Overlap.
Since the effective overlap $\rho$ depends only on the projection matrix $M$ and not on the parameter $\theta$, we measure $\rho$ without performing actual optimization. Instead, we randomly sample $M$ 1000 times for each method and compute $\rho$ accordingly.
The projection matrices $M$ are generated as follows:

### v2

@@ Appendix D Detailed Description of MeZO-BCD
This section provides a comprehensive description of our proposed method, MeZO-BCD, including its core design choices, the rationale for each, empirical comparisons, and algorithmic details. We focus on two critical dimensions: (1) how to define blocks for block-wise updates, and (2) how to select the active block at each iteration. For both, we describe the main alternatives considered, summarize supporting experiments, and justify our final recommendations.
D.1 Defining Blocks in MeZO-BCD
We primarily consider decoder-only Transformer architectures, which underlie most modern large language models (LLMs) (Zhang et al., 2022a; Touvron et al., 2023). These consist of an embedding layer (mapping input tokens to vectors; (embedding), a sequence of $L$ decoder layers (each with self-attention, feed-forward networks, and normalization layers), and a language model head (lm_head) projecting to token logits.
We evaluate three natural block partitioning strategies:
  • Layer-wise: The embedding, the lm_head, and each decoder layer are treated as individual blocks.
  • Linear-wise: The embedding, the lm_head, and each linear layer within every decoder block are treated as separate blocks.
  • Multi-layer: The embedding, the lm_head, and each block formed by grouping every two consecutive decoder layers together.
The layer-wise scheme aligns with functional modularity: decoder layers represent the principal processing units in LLMs, as recognized by both architectural conventions and empirical studies (Ju et al., 2024; Langedijk et al., 2024; Skean et al., 2025; Zhang et al., 2024d). This definition maintains interpretability and matches the typical tensor organization of model parameters. The linear-wise strategy increases block granularity by splitting each decoder layer into its constituent linear submodules, potentially enabling finer updates but at the cost of increased fragmentation. The multi-layer scheme, motivated by the sequential nature of hidden state propagation in transformers, merges every two consecutive decoder layers into a single block, reducing the number of blocks and further coarsening the update granularity.
It is important to note that defining blocks as arbitrary sub-tensors within a weight matrix (e.g., slicing a linear layer’s weight matrix) may appear to increase flexibility. However, it is impractical in modern deep learning frameworks since the entire tensor must still be loaded into memory for perturbation and update, negating any gain in memory transfer or efficiency.
Empirical Insights.
[…]
Table 3: Test accuracy of MeZO-BCD with different block partitioning schemes on OPT-1.3B. Results are the best accuracy across 20K steps, where test accuracies are evaluated every 4K steps.
Block Definition SST-2 RTE WIC
Linear-wise 88.1 57.4 58.2
Layer-wise 92.3 65.0 60.2
Multi-layer (2-layer) 91.2 65.0 58.2
Table 3 presents the results, showing that the linear-wise strategy consistently underperforms, even after hyperparameter tuning, whereas both layer-wise and multi-layer deliver similar results given appropriate learning rates. While smaller blocks reduce memory usage per step, excessively fine granularity can harm convergence. The layer-wise scheme offers the best trade-off between empirical accuracy and practical overhead, and is therefore adopted as the default in MeZO-BCD.
D.2 Block Selection Strategies in MeZO-BCD
Our theoretical analysis in the main text assumes that the active block is chosen i.i.d. uniformly at random at each step. However, in practical training, this can cause short-term “block starvation” periods where some blocks are updated much more or less frequently than others, which may hinder convergence stability.
To mitigate this, while preserving the marginal selection probability of $1/N$ for each block, we evaluate four block selection strategies:
  • Ascending order: $1\rightarrow 2\rightarrow\ldots\rightarrow N\rightarrow 1\rightarrow\ldots$
  • Descending order: $N\rightarrow N-1\rightarrow\ldots\rightarrow 1\rightarrow N\rightarrow\ldots$
  • Flip-flop order: $1\rightarrow 2\rightarrow\ldots\rightarrow N\rightarrow N-1\rightarrow\ldots\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow\ldots$
  • Cyclic random permutation: At the start of every $N$ steps, a random permutation of the $N$ blocks is generated and followed for the subsequent $N$ updates.
Here, a lower block index corresponds to layers closer to the input, and a higher index to layers closer to the output. While all strategies yield the same marginal probability of updating each block, they differ in their local ordering and thus in their short-term update patterns.
Empirical Insights.
To isolate the effect of block selection strategy, we conduct experiments with OPT-1.3B on SST-2, RTE, and WIC, fixing the block definition to layer-wise and searching over learning rates $\{5\times 10^{-6},1\times 10^{-5}\}$. The experimental protocol mirrors that in Table 3, ensuring comparability.
Table 4 presents the results. The flip-flop and cyclic random permutation strategies exhibit robust and consistently strong performance across all tasks, whereas the ascending and descending orderings show larger variance and less favorable results, particularly on RTE and WIC. Based on these findings, we adopt flip-flop and cyclic random permutation as our default block selection strategies, treating the choice between them as a discrete design parameter subject to tuning.
Table 4: Test accuracy of MeZO-BCD with different block selection strategies on OPT-1.3B. Results are the best accuracy across 20K steps, where test accuracies are evaluated every 4K steps
Block Selection SST-2 RTE WIC
Ascending order 91.9 61.7 58.0
Descending order 92.8 59.6 58.5
Flip-flop order 92.7 62.9 58.2
Cyclic random permutation 92.3 65.0 60.2
D.3 Theoretical Memory Consumption Analysis
To rigorously compare the memory consumption efficiency of zeroth-order methods, we analyze the theoretical peak memory requirements during training, as summarized in Table 5. Here, “peak” refers to the maximum instantaneous memory usage (e.g., VRAM) at any point during optimization, including temporary allocations for perturbations, but excluding activations or any other intermediate byproducts such as temporary forward buffers, dataloader prefetch, or I/O buffer.
Let the model consist of $L$ layers, with $\mathbf{W}_{\ell}\in\mathbb{R}^{m_{\ell}\times n_{\ell}}$ denoting the weight matrix of the $\ell$-th layer and $r$ the low-rank perturbation rank. For simplicity, we denote $|\mathbf{W}_{\ell}|\coloneqq m_{\ell}n_{\ell}$ by the total dimension of $\mathbf{W}_{\ell}$. Also, we refer to held perturbation as a segment of memory that is not continuously resident in the system but is instead allocated transiently in response to a Gaussian perturbation.
Table 5: Theoretical peak memory requirements (in number of parameters) during training, excluding activations or any other intermediate byproducts such as temporary forward buffers, dataloader prefetch, or I/O buffer. The max operator reflects instantaneous memory usage due to temporarily held perturbations.
Method Peak Memory Requirement
MeZO (Malladi et al., 2023) $\underbrace{\sum_{\ell=1}^{L}|\mathbf{W}_{\ell}|}_{\text{model weights}}+\underbrace{\max_{\ell}|\mathbf{W}_{\ell}|}_{\text{held perturbation}}$
SparseMeZO (Liu et al., 2024b) $\underbrace{\sum_{\ell=1}^{L}|\mathbf{W}_{\ell}|}_{\text{model weights}}+\underbrace{\max_{\ell}|\mathbf{W}_{\ell}|}_{\text{held perturbation}}+\underbrace{\max_{\ell}|\mathbf{W}_{\ell}|}_{\text{sparse mask}}$
LoZO (Chen et al., 2024) $\underbrace{\sum_{\ell=1}^{L}|\mathbf{W}_{\ell}|}_{\text{model weights}}+\underbrace{\max_{\ell}m_{\ell}r}_{\text{instant low-rank perturbation}}+\underbrace{\sum_{\ell=1}^{L}n_{\ell}r}_{\text{persistent low-rank factor}}$
MeZO-BCD (Ours) $\underbrace{\sum_{\ell=1}^{L}|\mathbf{W}_{\ell}|}_{\text{model weights}}+\underbrace{\max_{\ell\in B}|\mathbf{W}_{\ell}|}_{\text{held perturbation}}$
All methods must store the full set of model parameters ($\sum_{\ell=1}^{L}|\mathbf{W}_{\ell}|$) in memory throughout training. The additional held perturbation term represents the transient memory usage needed to apply weight perturbations within each training step. For MeZO (Malladi et al., 2023) and SparseMeZO (Liu et al., 2024b), this extra allocation never exceeds the size of a single layer’s weights at any moment, as perturbations are generated, used, and discarded immediately. SparseMeZO also incurs overhead from an additional mask of the same size. In contrast, LoZO (Chen et al., 2024)’s low-rank scheme involves both instant memory for sampled low-rank factors and persistent memory for factors reused across steps. MeZO-BCD, on the other hand, requires perturbing only the currently active block, so its peak auxiliary memory is, by construction, at most the largest weight size within any block $B$.
Importantly, although MeZO-BCD perturbs only one block at a time, every block is eventually selected during training. Therefore, the peak memory usage still corresponds to the largest block, and the effective bound matches that of MeZO. As a result, the true peak is determined by the largest layer across all blocks, making the effective memory bound identical to that of MeZO. Therefore, the use of the $\max$ operator in the analysis reflects that, while only one block is perturbed at a time, the overall peak is dictated by the largest such allocation encountered over all steps. Consequently, MeZO-BCD (like MeZO) achieves the theoretical minimum peak memory for zeroth-order optimization, essentially matching inference-time requirements, whereas alternative methods may require substantially more memory. This guarantees that MeZO-BCD’s temporary perturbations never incur a memory usage exceeding the size of the largest individual block or layer.
Remark on the experimental results in Table 1.
On SST-2, where activation buffers are relatively small, the reduction in auxiliary perturbation memory afforded by MeZO-BCD is large enough to lower the overall peak VRAM usage slightly, hence the small differences observed between methods on this dataset. By contrast, BoolQ and MultiRC use much longer input sequences, so their activation tensors dominate the GPU footprint. In those cases, the extra memory for perturbations or masks becomes negligible, and all methods converge to essentially the same peak memory.
Pseudocode for the full MeZO-BCD procedure is provided in Algorithm 1. We conclude this section with the following remark.
Why BCD is Particularly Effective in Zeroth-Order Optimization.
It is worth emphasizing why BCD, despite its classical status in first-order (FO) optimization, is particularly advantageous in the zeroth-order (ZO) context. In FO optimization, updating a single block still incurs the cost of computing and storing gradients for all downstream layers via backpropagation, resulting in substantial computational and memory overhead (especially for blocks closer to the input). In contrast, ZO optimization leverages only forward evaluations, so updating a block affects solely its own parameters with no added cost for the rest of the network. As a result, ZO-BCD achieves genuine reductions in memory transfer and computational overhead per step, avoiding the bottlenecks that limit the efficiency of BCD in first-order regimes. This distinction underlies the scalability and efficiency of MeZO-BCD for large-scale LLM fine-tuning.
Algorithm 1 MeZO-BCD
@@ D.3 Theoretical Memory Consumption Analysis
0:  block partitioning $\{B_{i}\}_{i=1}^{N}$, block ordering order, parameters $\mathbf{\theta}\in\mathbb{R}^{d}$ is partitioned into $[\theta_{B_{1}},\ldots,\theta_{B_{N}}]$, loss function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule $\{\eta_{t}\}_{t=1}^{T}$
@@
     Update current active block index $j\leftarrow\mathrm{\texttt{UpdateBlockIdx}}(\texttt{order},t,N)$
@@
Algorithm 2 PerturbParameters($\mathbf{\theta},\mu,s$)
0:  parameters $\mathbf{\theta}\in\mathbb{R}^{n}$, perturbation scale $\mu$, random seed $s$
0:
@@
Algorithm 3 UpdateBlockIdx(order, $t,N$)
0:  block ordering order, current time step $t$, number of blocks $N$
0:
  if order is ascending then
     $i\leftarrow((t-1)\bmod N)+1$
  else if order is descending then
     $i\leftarrow N-((t-1)\bmod N)$
  else if order is flip-flop then
     $i\leftarrow N-|((t-1)\bmod(2N-2))-(N-1)|$
  else if order is random then
     if $t\bmod N=0$ then
        Reset the random permutation $\pi\leftarrow\texttt{Shuffle}([N])$
     end if
     $i\leftarrow\pi[(t\bmod N)]$
@@ E.1 Randomized Quadratic Minimization
For the experiments in Section 3.4, the perturbation scale is globally set to $\mu=10^{-4}$ unless stated otherwise. The Hessian-like matrix $\mathbf{H}$ is generated as follows. The matrix is constructed as a block-diagonal structure, where each block is a randomly generated positive semi-definite matrix. Specifically, the $d\times d$ matrix $\mathbf{H}$ is divided into $B$ non-overlapping blocks, each of size $\frac{d}{B}\times\frac{d}{B}$. For each block, a random $\frac{d}{B}\times\frac{d}{B}$ matrix is sampled, and an orthogonal matrix $\mathbf{Q}_{i}$ is obtained via QR decomposition. Given a predefined maximum eigenvalue $\lambda_{\max}^{(i)}$ for each block, the diagonal matrix $\mathbf{\Lambda}^{(i)}$ is constructed as:
$\mathbf{\Lambda}^{(i)}=\mathrm{diag}(\underbrace{\lambda_{\max}^{(i)},\ldots,0.1\times\lambda_{\max}^{(i)}}_{r},\underbrace{0,\ldots,0}_{\frac{d}{B}-r}),$
@@
$\mathbf{H}_{i}=\mathbf{Q}_{i}\mathbf{\Lambda}^{(i)}\mathbf{Q}_{i}^{\top}.$
@@
This process ensures that $\mathbf{H}$ maintains a block-diagonal structure while preserving the spectral properties within each block. The following function constructs $\mathbf{H}$ according to the above procedure:
@@ E.1.1 Configuration for Figure 1
To examine the influence of the subspace alignment $\rho$, we use a dense Hessian matrix (i.e., $B=1$). The subspace alignment $\rho$ is controlled by introducing a parameter $\gamma\in[0,1]$, which determines the proportion of Hessian eigenvectors incorporated into the low-rank projection matrix $\mathbf{M}$. Specifically, $\lceil s\gamma\rceil$ eigenvectors (where $s=\mathrm{srank}(\mathbf{M})$) are randomly selected, while the remaining vectors are randomly generated and orthogonalized against the selected ones.
Formally, let the Hessian matrix $\mathbf{H}$ have rank $r$, and let $s$ denote the stable rank of $\mathbf{M}$. Denote the eigenvectors of $\mathbf{H}$ corresponding to nonzero eigenvalues as $\{\mathbf{q_{1}},\mathbf{q_{2}},\dots,\mathbf{q_{r}}\}$. To construct $\mathbf{M}$ with a controlled alignment $\rho$:
@@
    Randomly selected $s\gamma$ eigenvectors are set as columns of $\mathbf{M}_{1}$ (assuming $s\gamma$ is an integer for simplicity).
@@
    A random $d\times(1-s\gamma)$ matrix $\mathbf{R}$ is generated.
@@
    The QR decomposition is applied to $\mathbf{R}-\mathbf{M}_{1}\mathbf{M}_{1}^{\top}\mathbf{R}$ to obtain an orthogonal matrix $\mathbf{M}_{2}$ of size $d\times(1-s\gamma)$.
@@
    The final projection matrix is constructed as $\mathbf{M}=[\mathbf{M}_{1},\mathbf{M}_{2}]$.
@@
For this experiment, we set $d=256$, the Hessian rank to 64, and the stable rank $s$ of $\mathbf{M}$ to 64. This ensures that when $\gamma=1.0$, the projection matrix fully captures the Hessian eigenvectors. We use a learning rate of $\eta=10^{-3}$, a maximum of 1000 steps, and test five values of $\gamma$:
@@
Experimental Setup for the Middle Panel of Figure 1.
@@
  • The target loss is defined as the minimum loss achieved during training under $\gamma=0.0$.
@@
E.1.2 Configuration for the Right Panel of Figure 1
To evaluate the subspace alignment $\rho$ across different methods, we conduct randomized quadratic minimization experiments using a block-diagonal Hessian matrix with heterogeneous eigenspectra. To ensure a broader range of $\mathrm{srank}(\mathbf{M})$ values, the problem dimension is set to $d=1024$, with the Hessian consisting of 16 diagonal blocks, each of rank 16. To introduce heterogeneity in the eigenspectrum across blocks, the eigenvalues in each block are set by randomly choosing a reference from $\{10,40,70,100\}$ and sampling each eigenvalue as an integer within ±$\pm 2$ of it.
Measuring subspace alignment.
Since the subspace alignment $\rho$ depends only on the projection matrix $\mathbf{M}$ and not on the parameter $\theta$, we measure $\rho$ without performing actual optimization. Instead, we randomly sample $\mathbf{M}$ 1000 times for each method and compute $\rho$ accordingly.
The projection matrices $\mathbf{M}$ are generated as follows: